OpenStudy (anonymous):

Help. medal and fan,im no good at this 1/4 of the participants in a 5K race are children, and 2/7 of those are under 10 years old. What fraction of all of the participants are children under 10 years old? Select one: a. 1/14 b. 15/28 c. 13/14 d. 13/28

OpenStudy (anonymous):

\[\frac{ 1 }{4 }*\frac{ 2 }{ 7 }=\frac{ 1 }{ 14 }\]

OpenStudy (anonymous):

thanks man(:

OpenStudy (anonymous):

yw